[Brahms-l] Shift report 20040328 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Mar 28 2004 - 07:45:40 EST
------------------------------------------------------------
Date: Sun Mar 28 07:45:40 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 100 - 100

Shift Summary:
------------------------------------------------------------
1.  The machine has been operating very well.

2. I completed the data sets for MRS 45deg 6kG A, and FS 3deg 1/2A.  I also took four pedestal runs.

  Since we don't have a run plan written out, I followed the plan presented by Flemming last Thursday.  I moved the FS to 6deg, and I'm taking data at 1/4 field A pol. right now.

3.  I can not connect to the TOFW and CALB mainframes through BRAHVO, but I haven't noticed any problems with the detectors in the supermon plots. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 28 07:46:05 2004

This archive was generated by hypermail 2.1.8 : Sun Mar 28 2004 - 07:46:29 EST